OpenStudy (anonymous):

ok guys last one of the last one of the last one of the last........;) 105=x(x+2)(3) solve for x

OpenStudy (anonymous):

105=3x^2+6x, 3x^2+6x-105=0, 3(x^2+2x-35)=0, 3(x+7)(x-5)=0, x=-7,5 Do you understand?

OpenStudy (anonymous):

yesssssssssssssssssssssssssssssssssssssssssssss thanx thanx thanxxxxxxx